OpenStudy (anonymous):

Please Help Find the probability of at most one tail appearing on two tosses of a coin.

OpenStudy (anonymous):

wouldn't it just be a 50/50 chance since there are only two sides to a coin

OpenStudy (anonymous):

at most one tail = not two tails P(2 tails)=1/2*1/2=1/4 P(not 2 tails)=1-P(2 tails)=1-1/4

OpenStudy (anonymous):

The answer is correct .75. Thank You Very Much. I really appreciate it =)